# Nuisance-free Automatic Ground Collision Avoidance System Design: Merging Exponential-CBF and Adaptive Sliding Manifolds

### v1

@@ II-B Control Barrier Functions
where $\eta_{b}(\bm{x})=[b(\bm{x}),\mathcal{L}_{f}b(\bm{x}),\dots,\mathcal{L}^{\delta-1}_{f}b(\bm{x})]^{T}$ represents the Lie derivative vector of $b(\bm{x})$, and $\kappa=[k_{0},k_{1},\dots,k_{\delta-1}]$ denotes the coefficient gain vector associated with $\eta_{b}(\bm{x})$.
The components of $\kappa$ can be effectively determined using linear control techniques, such as pole placement [28, 29].
II-C Flight Dynamics Model
@@ II-C1 Equations of motion
$\begin{split}&\begin{cases}\dot{u}=\sum F_{x}/m+rv-qw\\ \dot{v}=\sum F_{y}/m+pw-ru\\ \dot{w}=\sum F_{z}/m+qu-pv\end{cases}\\ &\begin{cases}\dot{p}=qr(I_{\text{yy}}-I_{\text{zz}})/I_{\text{xx}}+(\dot{r}+pq)I_{\text{xz}}/I_{\text{xx}}+\sum L/I_{\text{xx}}\\ \dot{q}=pr(I_{\text{zz}}-I_{\text{xx}})/I_{\text{yy}}+(r^{2}-p^{2})I_{\text{xz}}/I_{\text{yy}}+\sum M/I_{\text{yy}}\\ \dot{r}=pq(I_{\text{xx}}-I_{\text{yy}})/I_{\text{zz}}+(\dot{p}-qr)I_{\text{xz}}/I_{\text{zz}}+\sum N/I_{\text{zz}}\end{cases}\\ &\begin{cases}\dot{x}_{\text{E}}=uc_{\theta}c_{\psi}+v(s_{\phi}s_{\theta}c_{\psi}-c_{\phi}s_{\psi})+w(c_{\phi}s_{\theta}c_{\psi}+s_{\phi}s_{\psi})\\ \dot{y}_{\text{E}}=uc_{\theta}c_{\psi}+v(s_{\phi}s_{\theta}c_{\psi}+c_{\phi}c_{\psi})+w(c_{\phi}s_{\theta}s_{\psi}-s_{\phi}s_{\psi})\\ \dot{z}_{\text{E}}=-us_{\theta}+vs_{\phi}c_{\theta}+wc_{\phi}c_{\theta}\end{cases}\\ &\begin{cases}\dot{\phi}=p+t_{\theta}(qs_{\phi}+rc_{\phi})\\ \dot{\theta}=qc_{\phi}-rs_{\phi}\\ \dot{\psi}=(qs_{\phi}+rc_{\phi})/c_{\theta}\end{cases}\\ \end{split}$ (6)
where $u$, $v$, and $w$ represent the body velocity components, while $p$, $q$, and $r$ are the body angular rate components. The navigational position components are denoted as $x_{E}$, $y_{E}$, and $z_{E}$, and $\phi$, $\theta$, and $\psi$ represent the Euler angles. The force components acting on the body frame are $F_{x}$, $F_{y}$, and $F_{z}$, and $L$, $M$, and $N$ correspond to the roll, pitch, and yaw moments. Finally, $m$ denotes the mass of the aircraft, and $I_{\text{xx}}$, $I_{\text{yy}}$, $I_{\text{zz}}$, and $I_{\text{xz}}$ are the moment of inertia components of the aircraft.
II-C2 Aerodynamics & Actuators
The over-actuated aerodynamic model is based on the methodology presented in [30, 31], where aerodynamic coefficients are expressed as polynomial functions of the relevant states and control surface deflections. Comprehensive details of this formulation are available in [30]. Additionally, the actuator dynamics are represented by a first-order system incorporating time constants, rate limits, and position saturation constraints, as described in [32]. Specifically, each control surface is characterized by a time constant of $0.0495s$. The rate saturation limits are $60^{\circ}/s$ for the horizontal tails, $80^{\circ}/s$ for the ailerons, and $120^{\circ}/s$ for the rudder. Similarly, the position saturation limits are ±$\pm 25^{\circ}$ for the horizontal tails, ±$\pm 21.5^{\circ}$ for the ailerons, and ±$\pm 30^{\circ}$ for the rudder.
II-D Flight Control Law Design
@@
II-D1 Control Augmentation System Design
@@
$\underbrace{\bm{\dot{\omega}}}_{\begin{subarray}{c}\dot{\bm{x}}\end{subarray}}=\underbrace{-J^{-1}(\bm{\omega}\times J\bm{\omega})}_{\begin{subarray}{c}\bm{f}(\bm{x})\end{subarray}}+\underbrace{J^{-1}\bar{q}_{\infty}S\begin{bmatrix}b&&\\ &\bar{c}&\\ &&b\\ \end{bmatrix}}_{\begin{subarray}{c}\bm{g}(\bm{x})\end{subarray}}\underbrace{\bm{\tau}}_{\begin{subarray}{c}\bm{u}\end{subarray}}$ (7)
where $\bm{\tau}$, $\bar{q}_{\infty}$, $S$, $b$, and $\bar{c}$ represent the aerodynamic moment coefficient vector, dynamic pressure, wing area, wing span, and mean aerodynamic chord, respectively. For a control-affine system as defined in Eq. (2), the NDI control law is expressed by Eq. (8).
$\bm{u}=\bm{g}(\bm{x})^{-1}[\bm{\nu}-\bm{f}(\bm{x})]$ (8)
@@
$\bm{\tau}_{c}=\Bigg{\{}J^{-1}\bar{q}_{\infty}S\begin{bmatrix}b&&\\ &\bar{c}&\\ &&b\\ \end{bmatrix}\Bigg{\}}^{-1}\Big{[}\bm{\dot{\omega}}_{c}+J^{-1}(\bm{\omega}\times J\bm{\omega})\Big{]}$ (9)
The virtual input $\bm{\dot{\omega}}_{c}$ is given by Eq. (10).
$\bm{\dot{\omega}}_{c}=\underbrace{\begin{bmatrix}K_{p}&&\\ &K_{q}&\\ &&K_{r}\\ \end{bmatrix}}_{\begin{subarray}{c}\overset{\Delta}{=}K\end{subarray}}\underbrace{\begin{bmatrix}p_{\text{pilot}}-p\\ q_{\text{pilot}}-q\\ r_{\text{pilot}}-r\\ \end{bmatrix}}_{\begin{subarray}{c}\overset{\Delta}{=}e\end{subarray}}$ (10)
where $K\overset{\Delta}{=}\{K_{p},K_{q},K_{r}\}$ denotes the gain matrix for roll, pitch, and yaw channels, respectively. These formulations provide the required control moment coefficients in response to pilot commands $p_{\text{pilot}}$, $q_{\text{pilot}}$, and $r_{\text{pilot}}$. The control moment coefficients are subsequently transferred to the control allocation module, detailed in the next section.
II-D2 Control Allocation Design
@@
$\Delta\bm{\delta}_{c}=\Phi^{-1}(\bm{x}_{0},\bm{\delta}_{0})\Delta\bm{\tau}_{c}$ (11)
where $\Delta\bm{\delta}_{c}=\bm{\delta}_{c}-\bm{\delta}_{0}$ and $\Delta\bm{\tau}_{c}=\bm{\tau}_{c}-\bm{\tau}_{0}$, with the subscript “00” denoting the current state. The control effectivity matrix $\Phi\in\mathbb{R}^{3\times n}$, shown in Eq. (12), contains moment coefficient derivatives with respect to control surface deflections at the current state, where $n$ represents the number of control surfaces. The control surface deflections $\bm{\delta}\in\mathbb{R}^{5}$ correspond to the right and left horizontal tails, right and left ailerons, and the rudder.
$\Phi(\bm{x}_{0},\bm{\delta}_{0})=\begin{bmatrix}{\dfrac{\partial C_{l}}{\partial\delta_{1}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}&\cdots&{\dfrac{\partial C_{l}}{\partial\delta_{n}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}\\ {\dfrac{\partial C_{m}}{\partial\delta_{1}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}&\cdots&{\dfrac{\partial C_{m}}{\partial\delta_{n}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}\\ {\dfrac{\partial C_{n}}{\partial\delta_{1}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}&\cdots&{\dfrac{\partial C_{n}}{\partial\delta_{n}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}\\ \end{bmatrix}$ (12)
@@
$\bm{\delta}_{c}=\Phi^{\ddagger}\Delta\bm{\tau}_{c}+\bm{\delta}_{0}$ (13)
@@
III Ground Collision Avoidance System Design
@@
$b(h)=h-(h_{\text{buff}}+h_{\text{DTED}})$ (14)
It is obvious that $b(h)>0,\forall h\in\mathbb{R}_{>(h_{\text{buff}}+h_{\text{DTED}})}$, and $b(h)=0,h=h_{\text{buff}}+h_{\text{DTED}}$. Thus, the time derivative of the barrier function is in Eq. (15).
$\dot{b}(h)=\dot{h}=\mathcal{L}_{f}b(h)$ (15)
However, in the first time derivative, desired inputs are not observable, thereby the second time derivative of the barrier function is presented in Eq. (16).
$\ddot{b}(h)=\ddot{h}=f(h)+g(h)u=\mathcal{L}^{2}_{f}b(h)+\mathcal{L}_{g}\mathcal{L}_{f}b(h)u$ (16)
The roll and pitch rates can be explicitly obtained by expanding Eq. (16); however, for simplicity, the following assumption in III.1 is both convenient and consistent.
Assumption III.1.
Assume that the bank angle ($\phi$) and sideslip angle ($\beta$) are both $0^{\circ}$, as wing-level and symmetric flight is expected during the ground collision avoidance maneuver.
@@
$\dot{h}=V_{T}s_{\gamma}$ (17)
where $V_{T}=||\bm{V}||_{2}$, note that $\bm{V}=[u\hskip 4.26773ptv\hskip 4.26773ptw]^{T}$, is the true airspeed, and $\gamma=\theta-\alpha$ due to the wings-level symmetric flight. Consequently, the altitude dynamics are reduced to dependencies on true airspeed, pitch angle, and angle of attack. At this stage, these dynamics should be decomposed to isolate the pitch rate. Based on this rationale, the second time derivative of the barrier function to observe the pitch rate is given by Eq. (18).
@@
$=\big{[}f(V_{T})+g(V_{T})q\big{]}s_{(\theta-\alpha)}+\cdots$
$\cdots+V_{T}c_{(\theta-\alpha)}\big{[}f(\theta)+g(\theta)q-f(\alpha)-g(\alpha)q\big{]}$
@@
$\ddot{h}$ $=\underbrace{f(V_{T})+V_{T}c_{(\theta-\alpha)}\big{[}f(\theta)-f(\alpha)\big{]}}_{\begin{subarray}{c}f(h)\end{subarray}}+\cdots$ (19)
$\cdots+\underbrace{\Big{(}g(V_{T})s_{(\theta-\alpha)}+V_{T}c_{(\theta-\alpha)}\big{[}g(\theta)-g(\alpha)\big{]}\Big{)}}_{\begin{subarray}{c}g(h)\end{subarray}}q$
At this point, the dynamics of true airspeed, pitch angle, and angle of attack should be defined; however, the detailed derivation of these dynamics is not included in this study, as it is readily available in the flight dynamics and control literature, e.g. [35, 36]. The true airspeed dynamics are given by Eq. (20).
$\dot{V}_{T}$ $=\dfrac{1}{m}\big{[}-Dc_{\beta}+Cs_{\beta}+Tc_{\alpha}c_{\beta}\cdots$ (20)
$\cdots-mg\big{(}s_{\theta}c_{\alpha}c_{\beta}-c_{\theta}s_{\phi}s_{\beta}-c_{\theta}c_{\phi}s_{\alpha}c_{\beta}\big{)}\big{]}$
@@
$\dfrac{-Dc_{\beta}}{m}=\dfrac{-\bar{q}_{\infty}SC_{D_{q}}\dfrac{q\bar{c}}{2V_{T}}c_{\beta}}{m}+\dfrac{-\bar{q}_{\infty}SC_{D_{\text{aux}}}c_{\beta}}{m}$ (21)
This decomposition should be done to reveal hidden pitch rate contributions within the aerodynamic coefficients, where $C_{D_{\text{aux}}}$ represents remaining lift coefficient terms excluding $C_{D_{q}}$. For the simplicity, the assumption in III.1 also leads to the following assumption described in III.2.
Assumption III.2.
@@
$f(V_{T})$ $=\frac{Tc_{\alpha}c_{\beta}-mg\big{(}s_{\theta}c_{\alpha}c_{\beta}-c_{\theta}s_{\phi}s_{\beta}-c_{\theta}c_{\phi}s_{\alpha}c_{\beta}\big{)}}{m}\cdots$ (22)
$\cdots-\dfrac{\bar{q}_{\infty}SC_{D_{\text{aux}}}c_{\beta}}{m}$
$g(V_{T})$ $=-\dfrac{\bar{q}_{\infty}SC_{D_{q}}\dfrac{\bar{c}}{2V_{T}}c_{\beta}}{m}$
@@
$\dot{\alpha}$ $=\dfrac{-L}{mV_{T}c_{\beta}}+\dfrac{mg\big{(}c_{\theta}c_{\phi}c_{\alpha}+s_{\theta}s_{\alpha}\big{)}-T}{mV_{T}c_{\beta}}+q\cdots$ (25)
$\cdots-t_{\beta}\big{(}pc_{\alpha}+rs_{\alpha}\big{)}$
@@
$\dfrac{-L}{mV_{T}c_{\beta}}=\dfrac{-\bar{q}_{\infty}SC_{L_{q}}\dfrac{q\bar{c}}{2V_{T}}}{mV_{T}c_{\beta}}+\dfrac{-\bar{q}_{\infty}SC_{L_{\text{aux}}}}{mV_{T}c_{\beta}}$ (26)
@@
$f(\alpha)$ $=\frac{mg\big{(}c_{\theta}c_{\phi}c_{\alpha}+s_{\theta}s_{\alpha}\big{)}-T}{mV_{T}c_{\beta}}\cdots$ (27)
$\cdots-t_{\beta}\big{(}pc_{\alpha}+rs_{\alpha}\big{)}-\frac{\bar{q}_{\infty}SC_{L_{\text{aux}}}}{mV_{T}c_{\beta}}$
$g(\alpha)$ $=1-\dfrac{\bar{q}_{\infty}SC_{L_{q}}\frac{\bar{c}}{2V_{T}}}{mV_{T}c_{\beta}}$
@@
$\eta_{b}(h)=\begin{bmatrix}b(h)\\ \mathcal{L}_{f}b(h)\\ \end{bmatrix}=\begin{bmatrix}h-(h_{\text{buff}}+h_{\text{DTED}})\\ \dot{h}\\ \end{bmatrix}$ (28)
@@
$\underbrace{f(h)+g(h)q_{\text{GCAS}}}_{\begin{subarray}{c}\mathcal{L}^{2}_{f}b(h)+\mathcal{L}_{g}\mathcal{L}_{f}b(h)q\end{subarray}}+\kappa\eta_{b}(h)\geq 0$ (29)
@@
$\min_{q_{\text{GCAS}}\in\mathbb{R}}$ $\dfrac{1}{2}\big{(}q_{\text{GCAS}}-q_{\text{pilot}}\big{)}^{2}$ (30)
@@
$q_{\text{min}}\leq q_{\text{GCAS}}\leq q_{\text{max}}$
The constructed formulation enables the Auto-GCAS to generate a pitch rate command ($q_{\text{GCAS}}$) that closely follows the pitch rate command of pilot ($q_{\text{pilot}}$) while adhering to the constraint for ground collision avoidance. Additionally, the generated pitch rate command, $q_{\text{GCAS}}$, must remain within the interval $[q_{\text{min}}\hskip 4.26773ptq_{\text{max}}]$, considering the admissible and allowable pitch rate limits depending on the aircraft. However, at this stage, the nuisance-free criterion has not been addressed, and the question of whether it is possible to generate nuisance-free commands by adjusting the gains of $\kappa$, i.e., $k_{1}$ and $k_{2}$, will be elaborated in the proceeding section.
III-A Nuisance-free Intervention Design

### v2

@@ 2.2 Control Barrier Functions
where $\eta_{b}(\bm{x})=[b(\bm{x}),\mathcal{L}_{f}b(\bm{x}),...,\mathcal{L}_{f}^{\delta-1}b(\bm{x})]^{\mathrm{T}}$ is the Lie derivative vector of $b(\bm{x})$ and $\kappa=[k_{0},k_{1},\dots,k_{\delta-1}]$ denotes the coefficient gain vector associated with $\eta_{b}(\bm{x})$.
2.3 Flight Dynamics Model
@@ 2.3.1 Equations of motion
[…]
where $u\hskip 4.26773ptv\hskip 4.26773ptw]^{\mathrm{T}}\in\mathbb{R}^{3}$ represent the body velocity components, while $p\hskip 4.26773ptq\hskip 4.26773ptr]^{\mathrm{T}}\in\mathbb{R}^{3}$ are the body angular rate components. The navigational position components are denoted as $x_{\text{E}}\hskip 4.26773pty_{\text{E}}\hskip 4.26773ptz_{\text{E}}]^{\mathrm{T}}\in\mathbb{R}^{3}$, and $\phi\hskip 4.26773pt\theta\hskip 4.26773pt\psi]^{\mathrm{T}}\in\mathbb{R}^{3}$ represent the Euler angles. The force components acting on the body frame are $F_{x}\hskip 4.26773ptF_{y}\hskip 4.26773ptF_{z}]^{\mathrm{T}}\in\mathbb{R}^{3}$, and $L\hskip 4.26773ptM\hskip 4.26773ptN]^{\mathrm{T}}\in\mathbb{R}^{3}$ correspond to the roll, pitch, and yaw moments. Finally, $m$ denotes the mass of the aircraft, and $I_{\text{xx}}$, $I_{\text{yy}}$, $I_{\text{zz}}$, and $I_{\text{xz}}$ are the moment of inertia components of inertia tensor $J\in\mathbb{R}^{3\times 3}$ of the aircraft.
2.3.2 Aerodynamics & Actuators
The over-actuated aerodynamic model is based on the methodology presented in [27, 28], where aerodynamic coefficients are expressed as polynomial functions of the relevant states and control surface deflections. Comprehensive details of this formulation are available in [27]. Additionally, the actuator dynamics are represented by a first-order system incorporating time constants, rate limits, and position saturation constraints, as described in [29].
2.4 Flight Control Law Design
@@
2.4.1 Control Augmentation System Design
@@
$\underbrace{\bm{\dot{\omega}}}_{\begin{subarray}{c}\dot{\bm{x}}\end{subarray}}=\underbrace{-J^{-1}(\bm{\omega}\times J\bm{\omega})}_{\begin{subarray}{c}\bm{f}(\bm{x})\end{subarray}}+\underbrace{J^{-1}\bar{q}_{\infty}S\begin{bmatrix}b&&\\ &\bar{c}&\\ &&b\\ \end{bmatrix}}_{\begin{subarray}{c}\bm{g}(\bm{x})\end{subarray}}\underbrace{\bm{\tau}}_{\begin{subarray}{c}\bm{u}\end{subarray}},$ (8)
where $\bm{\tau}\in\mathbb{R}^{3}$, $\bar{q}_{\infty}\in\mathbb{R}$, $S\in\mathbb{R}$, $b\in\mathbb{R}$, and $\bar{c}\in\mathbb{R}$ represent the aerodynamic moment coefficient vector, dynamic pressure, wing area, wing span, and mean aerodynamic chord, respectively. For a control-affine system as defined in Eq. (2), the NDI control law is expressed by Eq. (9).
$\bm{u}=\bm{g}(\bm{x})^{-1}[\bm{\nu}-\bm{f}(\bm{x})],$ (9)
@@
$\bm{\tau}_{c}=\Bigg{\{}J^{-1}\bar{q}_{\infty}S\begin{bmatrix}b&&\\ &\bar{c}&\\ &&b\\ \end{bmatrix}\Bigg{\}}^{-1}\Big{[}\bm{\dot{\omega}}_{c}+J^{-1}(\bm{\omega}\times J\bm{\omega})\Big{]}.$ (10)
The virtual input $\bm{\dot{\omega}}_{c}\in\mathbb{R}^{3}$ is given by Eq. (11).
$\bm{\dot{\omega}}_{c}=\underbrace{\begin{bmatrix}K_{p}&&\\ &K_{q}&\\ &&K_{r}\\ \end{bmatrix}}_{\begin{subarray}{c}\overset{\Delta}{=}K\end{subarray}}\underbrace{\begin{bmatrix}p_{\text{pilot}}-p\\ q_{\text{pilot}}-q\\ r_{\text{pilot}}-r\\ \end{bmatrix}}_{\begin{subarray}{c}\overset{\Delta}{=}e\end{subarray}},$ (11)
where $K\overset{\Delta}{=}\{K_{p},K_{q},K_{r}\}$ denotes the gain matrix for roll, pitch, and yaw channels, respectively. These formulations provide the required control moment coefficients in response to pilot commands $p_{\text{pilot}}\in\mathbb{R}$, $q_{\text{pilot}}\in\mathbb{R}$, and $r_{\text{pilot}}\in\mathbb{R}$. The control moment coefficients are subsequently transferred to the control allocation module, detailed in the next section.
2.4.2 Control Allocation Design
@@
$\Delta\bm{\delta}_{c}=\Phi^{-1}(\bm{x}_{0},\bm{\delta}_{0})\Delta\bm{\tau}_{c},$ (12)
where $\Delta\bm{\delta}_{c}=\bm{\delta}_{c}-\bm{\delta}_{0}$ and $\Delta\bm{\tau}_{c}=\bm{\tau}_{c}-\bm{\tau}_{0}$, with the subscript “0” denoting the current state. The control effectivity matrix $\Phi\in\mathbb{R}^{3\times n}$, shown in Eq. (13), contains moment coefficient derivatives with respect to control surface deflections at the current state, where $n$ represents the number of control surfaces. The control surface deflections $\bm{\delta}\in\mathbb{R}^{5}$ correspond to the right and left horizontal tails, right and left ailerons, and the rudder.
$\Phi(\bm{x}_{0},\bm{\delta}_{0})=\begin{bmatrix}{\dfrac{\partial C_{l}}{\partial\delta_{1}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}&\cdots&{\dfrac{\partial C_{l}}{\partial\delta_{n}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}\\ {\dfrac{\partial C_{m}}{\partial\delta_{1}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}&\cdots&{\dfrac{\partial C_{m}}{\partial\delta_{n}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}\\ {\dfrac{\partial C_{n}}{\partial\delta_{1}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}&\cdots&{\dfrac{\partial C_{n}}{\partial\delta_{n}}}\Big{|}_{(\bm{x}_{0},\bm{\delta}_{0})}\\ \end{bmatrix}$. (13)
@@
$\bm{\delta}_{c}=\Phi^{\ddagger}\Delta\bm{\tau}_{c}+\bm{\delta}_{0}.$ (14)
@@
3 Ground Collision Avoidance System Design
@@
$b(h)=h-(h_{\text{buff}}+h_{\text{DTED}}).$ (15)
It is obvious that $b(h)>0,\forall h\in(h_{\text{buff}}+h_{\text{DTED}},\infty)$, and $b(h)=0\iff h=h_{\text{buff}}+h_{\text{DTED}}$. Thus, the time derivative of the barrier function is in Eq. (16).
$\dot{b}(h)=\dot{h}=\mathcal{L}_{f}b(h).$ (16)
However, in the first time derivative, desired input does not appear, thereby the second time derivative of the barrier function is presented in Eq. (17).
$\begin{split}\ddot{b}(h)=\ddot{h}&=f(h)+g(h)u\\ &=\mathcal{L}^{2}_{f}b(h)+\mathcal{L}_{g}\mathcal{L}_{f}b(h)u.\end{split}$ (17)
The roll and pitch rates can be explicitly obtained by expanding Eq. (17); however, for simplicity, the following assumption in 3.1 is both convenient and consistent.
Assumption 3.1.
We assume that the bank angle ($\phi$) and sideslip angle ($\beta$) are both $0^{\circ}$, as wing-level and symmetric flight is expected during the ground collision avoidance maneuver.
@@
$\dot{h}=V_{T}s_{\gamma},$ (18)
[…]
@@
$=\big{[}f(V_{T})+g(V_{T})q\big{]}s_{(\theta-\alpha)}+V_{T}c_{(\theta-\alpha)}\big{[}f(\theta)+g(\theta)q-f(\alpha)-g(\alpha)q\big{]}.$
@@
$\ddot{h}$ $=\underbrace{f(V_{T})+V_{T}c_{(\theta-\alpha)}\big{[}f(\theta)-f(\alpha)\big{]}}_{\begin{subarray}{c}f(h)\end{subarray}}+\underbrace{\Big{(}g(V_{T})s_{(\theta-\alpha)}+V_{T}c_{(\theta-\alpha)}\big{[}g(\theta)-g(\alpha)\big{]}\Big{)}}_{\begin{subarray}{c}g(h)\end{subarray}}q.$ (20)
At this point, the dynamics of true airspeed, pitch angle, and angle of attack should be defined; however, the detailed derivation of these dynamics is not included in this study, as it is readily available in the flight dynamics and control literature, e.g. [32, 33, 34]. The true airspeed dynamics are given by Eq. (21).
$\dot{V}_{T}$ $=\dfrac{1}{m}\big{[}-Dc_{\beta}+Cs_{\beta}+Tc_{\alpha}c_{\beta}-mg\big{(}s_{\theta}c_{\alpha}c_{\beta}-c_{\theta}s_{\phi}s_{\beta}-c_{\theta}c_{\phi}s_{\alpha}c_{\beta}\big{)}\big{]},$ (21)
@@
$\dfrac{-Dc_{\beta}}{m}=\dfrac{-\bar{q}_{\infty}SC_{D_{q}}\dfrac{q\bar{c}}{2V_{T}}c_{\beta}}{m}+\dfrac{-\bar{q}_{\infty}SC_{D_{\text{aux}}}c_{\beta}}{m}.$ (22)
This decomposition should be done to reveal hidden pitch rate contributions within the aerodynamic coefficients, where $C_{D_{\text{aux}}}$ represents remaining lift coefficient terms excluding $C_{D_{q}}$. For the simplicity, the assumption in 3.1 also leads to the following assumption described in 3.2.
Assumption 3.2.
@@
$f(V_{T})$ $=\frac{Tc_{\alpha}c_{\beta}-mg\big{(}s_{\theta}c_{\alpha}c_{\beta}-c_{\theta}s_{\phi}s_{\beta}-c_{\theta}c_{\phi}s_{\alpha}c_{\beta}\big{)}}{m}-\dfrac{\bar{q}_{\infty}SC_{D_{\text{aux}}}c_{\beta}}{m},$ (23)
$g(V_{T})$ $=-\dfrac{\bar{q}_{\infty}SC_{D_{q}}\dfrac{\bar{c}}{2V_{T}}c_{\beta}}{m}.$
@@
$\dot{\alpha}$ $=\dfrac{-L}{mV_{T}c_{\beta}}+\dfrac{mg\big{(}c_{\theta}c_{\phi}c_{\alpha}+s_{\theta}s_{\alpha}\big{)}-T}{mV_{T}c_{\beta}}+q-t_{\beta}\big{(}pc_{\alpha}+rs_{\alpha}\big{)},$ (26)
@@
$\dfrac{-L}{mV_{T}c_{\beta}}=\dfrac{-\bar{q}_{\infty}SC_{L_{q}}\dfrac{q\bar{c}}{2V_{T}}}{mV_{T}c_{\beta}}+\dfrac{-\bar{q}_{\infty}SC_{L_{\text{aux}}}}{mV_{T}c_{\beta}},$ (27)
@@
$f(\alpha)$ $=\frac{mg\big{(}c_{\theta}c_{\phi}c_{\alpha}+s_{\theta}s_{\alpha}\big{)}-T}{mV_{T}c_{\beta}}-t_{\beta}\big{(}pc_{\alpha}+rs_{\alpha}\big{)}-\frac{\bar{q}_{\infty}SC_{L_{\text{aux}}}}{mV_{T}c_{\beta}},$ (28)
$g(\alpha)$ $=1-\dfrac{\bar{q}_{\infty}SC_{L_{q}}\frac{\bar{c}}{2V_{T}}}{mV_{T}c_{\beta}}.$
@@
$\eta_{b}(h)=\begin{bmatrix}b(h)\\ \mathcal{L}_{f}b(h)\\ \end{bmatrix}=\begin{bmatrix}h-(h_{\text{buff}}+h_{\text{DTED}})\\ \dot{h}\\ \end{bmatrix}$. (29)
@@
$\underbrace{f(h)+g(h)q_{\text{GCAS}}}_{\begin{subarray}{c}\mathcal{L}^{2}_{f}b(h)+\mathcal{L}_{g}\mathcal{L}_{f}b(h)q\end{subarray}}+\kappa\eta_{b}(h)\geq 0,$ (30)
@@
$u^{\star}=\operatorname*{argmin}_{q_{\text{GCAS}}\in\mathbb{R}}$ $\dfrac{1}{2}\big{(}q_{\text{GCAS}}-q_{\text{pilot}}\big{)}^{2}$ (31)
@@
$q_{\text{min}}\leq q_{\text{GCAS}}\leq q_{\text{max}}.$
The constructed formulation enables the Auto-GCAS to generate a pitch rate command ($q_{\text{GCAS}}$) that closely follows the pitch rate command of pilot ($q_{\text{pilot}}$) while adhering to the constraint for ground collision avoidance. Additionally, the generated pitch rate command, $q_{\text{GCAS}}$, must remain within the interval $q_{\text{min}}\hskip 4.26773ptq_{\text{max}}]$, considering the admissible and allowable pitch rate limits depending on the aircraft. However, at this stage, the nuisance-free criterion has not been addressed, and the question of whether it is possible to generate nuisance-free commands by adjusting the gains of $\kappa$, i.e., $k_{1}$ and $k_{2}$, will be elaborated in the proceeding section.
3.1 Nuisance-free Intervention Design